OpenStudy (zasharra):

Find all solutions in the interval [0, 2π). 7 tan2x - 21 tan x = 0

OpenStudy (zasharra):

@mathmate

OpenStudy (zasharra):

@tkhunny

OpenStudy (irishboy123):

i'd say always rip it into sin and cos i reckon you have: \(7 \tan2x - 21 \tan x = 0\)

OpenStudy (zasharra):

I have the numbers but i need it in fraction form

OpenStudy (zasharra):

^2

OpenStudy (zasharra):

7 tan ^2x - 21 tan x=0

OpenStudy (tkhunny):

Treat it like a quadratic equation. Factor and solve.

OpenStudy (tkhunny):

No sine and cosine silliness. You should know where tan(x) = 0 and tan(x) = 3.

OpenStudy (irishboy123):

Mmmmm \(7 \dfrac{\sin 2x}{\cos 2x} - 21 \dfrac{\sin x}{\cos x} = 0\) \( \dfrac{\sin 2x}{\cos 2x} - 3 \dfrac{\sin x}{\cos x} = 0\) \( \dfrac{2 \sin x \cos x}{\cos 2x} - 3 \dfrac{\sin x}{\cos x} = 0\) nil solution is \(\sin x = 0\)....... \( \dfrac{2 \cos x}{\cos 2x} - 3 \dfrac{1}{\cos x} = 0\) \( \dfrac{2 \cos^2 x}{1} - 3 \dfrac{\cos 2x}{1} = 0\) Hang on a mo, its: \(7 \tan^2(x) - 21 \tan( x) = 0\) And not \(7 \tan(2x) - 21 \tan (x) = 0\) Grrrr